# MOSAIC: Modeling Social AI for Content Dissemination and Regulation in Multi-Agent Simulations

### v2

@@ Abstract
We present a novel, open-source social network simulation framework MOSAIC where generative language agents predict user behaviors such as liking, sharing, and flagging content. This simulation combines LLM agents with a directed social graph to analyze emergent deception behaviors and gain a better understanding of how users determine the veracity of online social content. By constructing user representations from diverse fine-grained actual user personas, our system enables multi-agent simulations that model content dissemination and engagement dynamics at scale. Within this framework, we evaluate three different content moderation strategies with simulated misinformation dissemination, and we find that they not only mitigate the spread of non-factual content but also increase user engagement. In addition, we analyze the trajectories of popular content in our simulations, and explore whether simulation agents’ articulated reasoning for their social interactions truly aligns with their collective engagement patterns. We open-source our simulation software to encourage further research within AI and social sciences: https://github.com/AnonymousPaperRepo/MOSAIC-submission
@@ 1 Introduction
In 2024, OpenAI reported that its platform was already being misused by covert influence operations to generate synthetic content diffused over social media OpenAI (2024). These internet manipulators exploit the fact that social networks have become a fundamental part of modern life, shaping public discourse, influencing political opinions, and facilitating the rapid spread of unverified human- and AI-generated content Aichner et al. (2021); Orben et al. (2022); Cinelli et al. (2021). While traditional social science methods such as surveys and observational studies have provided insights into human behavior, they often struggle to capture large-scale, emergent online interactions Yu et al. (2021); Lorig et al. (2021). Agent-based modeling (ABM) provides distinct advantages over survey methods in social science research since they can simulate dynamic interactions over time, and support examination of hypothetical or counterfactual scenarios with repeatable and controllable conditions Bonabeau (2002); Epstein (1999).
@@
By bridging social science observations, game-theoretic modeling Acemoglu et al. (2023) and LLM-driven modeling, our work demonstrates the potential of generative agent simulations as a tool for studying large-scale online behaviors.
@@ 2 MOSAIC Social Network Simulation
We build a simulated social network environment inspired by platforms like X, [^1] https://x.com/ allowing AI-driven users to interact, post, and share content. The simulation includes a basic user class with attributes such as username, posts, followers, following, and reposts, mimicking the structure of real-world social media platforms. The network itself is defined by the follower-following relationships, creating a web of user interactions, represented by a directed graph $G=(N,E)$ where $N$ represents the set of user nodes, i.e., $N=\{n_{1},n_{2},\dots,n_{k}\}$, where $n_{i}$ is a user in the network. $E\subseteq N\times N$ represents the set of directed edges, i.e., $E=\{(n_{i},n_{j})\mid n_{i}\text{ follows }n_{j}\}$. Each edge $(n_{i},n_{j})$ signifies that user $n_{i}$ follows user $n_{j}$.
@@ 2.1 Simulation Flow
At each time step, news content is introduced based on predefined parameters, with agents dynamically responding to their feeds. Agents can optionally generate posts according to their own interests. However during certain controlled experiments we configure them to only engage through reactions such as liking, sharing, commenting, or reporting misinformation. We describe a more general action space and more details of their decision-making process in Appendix G. The visibility of posts evolves based on engagement metrics, simulating algorithmic amplification effects. If fact-checking is enabled, agents incorporate moderation signals, e.g. they are prompted to pay more attention to potentially falsified content or misinformation. We discuss the content moderation simulation in more depth in Section 3. Throughout a simulation run, the system tracks key statistics, including content reach, user influence, and misinformation spread. At the end of each simulation run, a post-hoc analysis is conducted to assess content diffusion dynamics, user engagement metrics, influence distribution, and the impact of fact-checking interventions. We also keep track of various network properties such as centrality [^2] The degree to which an user is central to the network, having outsized or undersized influence. and triadic closures, [^3] A common social phenomenon where users with a mutual connection are more likely to connect to each other. and perform homophily [^4] Another common social phenomenon where similar users are more likely to connect to each other over dissimilar ones. analysis to examine clustering patterns in user engagement (see Appdendix A for an extended technical description of how we define and compute these network metrics). Our focus on these three properties is motivated by a large body of prior work in network theory and the social sciences Rapoport (1953); Granovetter (1973); McPherson et al. (2001); Abebe et al. (2022); Chang et al. (2025).
@@ 3.1 Setup
To collect factual news articles, we utilize a news aggregation API [^8] https://newsapi.org/ to retrieve articles published daily in legitimate sources from January 31 to February 28, 2025. The system queries the API for all available topics, prioritizing non-political popular articles in English. For each date in the range, we extract key information from the retrieved articles, including their title, description and main content. Using the NewsAPI, we scraped a total number of 2470 pieces of news articles from major media outlets.
@@
Our simulations involve agentic users interacting with news posts under four different fact-checking conditions: (1) No Fact-Checking, (2) Community-Based Fact-Checking, (3) Third-Party Fact-Checking with an independent LLM that uses its own parametric knowledge, and (4) Hybrid Fact-Checking, which integrates both community-based and third-party verification mechanisms. The simulations starts with 50 agents, and spans over 40 time steps, with agents making interaction decisions based on the perceived veracity of posts and the presence of content moderation. At each new step, we randomly introduce up to 2 more agents into the environment to simulate the regular user growth of the social media platform. We analyze both the overall engagement with posts and the effectiveness of fact-checking strategies in suppressing misinformation.
@@
The action space of agents varies across different fact-checking conditions, reflecting different levels of scrutiny and intervention in their social media interactions. In the no fact-checking setting, agents interact freely with the feed, engaging with posts based solely on their interests and beliefs. They can like, share, comment, or ignore posts, without any explicit instructions to assess the accuracy of the content. Under this setting, a post is considered “fact-checked" when it has at least one community note with a helpful rating. In the third-party fact-checking condition, the action space remains the same, but the environment implicitly assumes the presence of external fact-checkers who may influence the visibility or credibility of posts. However, the agents themselves do not perform any direct verification. In contrast, the community fact-checking setting expands the action space by allowing agents to add community notes to posts they deem misleading or in need of additional context, as well as rate existing community notes as either helpful or unhelpful. This introduces a participatory element, encouraging agents to contribute to a crowdsourced verification system. Finally, the hybrid fact-checking condition combines elements from both third-party and community-driven verification. Agents can engage with posts as in previous settings while also considering official fact-checks alongside community notes, contributing their own notes and rating those written by others. Across all conditions, agents must select from predefined valid actions, ensuring consistency in response formats. Additionally, when reasoning is enabled, agents are required to justify their interactions by providing a brief explanation for each chosen action, further enhancing the interpretability of their behavior.
@@ 3.2 False News Does Not Spread Faster than Real News with Simulation Agents
As shown in Fig. 4, factual news maintains consistently higher engagement levels than misinformation across all four settings. The gap is most pronounced under the Third-Party and Hybrid fact-checking conditions, where factual interactions climb steadily while misinfo remains low. Notably, in the No Fact-Checking scenario, false news still fails to gain a foothold, suggesting that these agents are inherently less likely to propagate unverified or misleading posts. We believe this behavior arises from two factors. First, the agents may rely on their internal confidence to guide sharing decisions, so low-confidence (often false) content is passed over more often. And second, commercial LLMs have been post-trained with safety and helpfulness objectives. This training bias would make our agents less inclined to produce or circulate uncertain or harmful content.
@@ 3.3 Content Moderation Improves Both Fact-Checking and Engagement
The steadily widening gaps between the third-party/hybrid approaches indicate that certain content moderation generates compounding benefits for factual engagement over time. Among the simulated strategies, fact-checking by third parties proves most effective at building a healthy information ecosystem—an effect that aligns with human studies showing increased trust in unflagged content Pennycook et al. (2020). All curves in Fig. 4 represent metrics averaged across our three LLM agents (DeepSeek-V3, GPT-4o, and Claude-3.7-Sonnet). Interaction counts including the number of shares, likes, comments, and notes, are aggregated at each time step to highlight overall platform dynamics under each moderation strategy. This high-level view illustrates consistent patterns in how content moderation shapes engagement. For a detailed, model-by-model breakdown and finer-grained metric analysis, please see Appendix J.
@@ 3.5 What Drives Users/Content Popularity?
Understanding what makes some content more popular is key for modeling engagement and intervention. In this standalone study, we simulate 161 agents over 4,249 posts of news and user-generated content, and let agents both react to and generate content based on personas, engagement cues, and memory-based decisions.
@@
Finally, we examine agents’ own reasoning traces to better understand engagement behaviors. Interestingly, agent reasoning traces show that sentiment and motivations do not always align with action types (e.g., positive sentiment in follows, likes, shares; quality or misinformation concerns in flags). As shown in Tab. 3, 13.5% of “flags” are coded as positive even though flag reasons are exclusively negative (information value, misinformation). The verbal reasoning for the “unfollow” actions is 20% positive though it is intuitively a negative reaction towards other users. “ignore” omits any neutral reasoning and a 77.8% positive rate in reasoning, which also makes their verbal explanations inconsistent with the actual action. Following these inconclusive analyses, we speculate that our recency-and-follow-based feed ranking creates preferential attachment: once followed, a user’s posts gain visibility, reinforcing their popularity. We provide further analyses in Appendix A.
@@ 5 Conclusion
Our findings are subject to several limitations, particularly in the scale of our experiments. First, the limited number of human participants, especially from minority demographic groups, restricts the statistical power of our conclusions. Expanding participant diversity would enable a more robust analysis of how alignment between real and simulated social interaction patterns varies across demographics. Second, our content moderation experiments were conducted at a relatively small scale, which may have constrained the emergence of complex behaviors. Running these experiments at a larger scale could uncover additional dynamics not captured in the present study. Finally, we observe a gap between agents’ explicit explanations for their actions and the collective reaction patterns that emerge in the system. The root causes of this misalignment remain unclear and warrant further investigation, potentially involving a deeper analysis of agent modeling assumptions or social influence mechanisms.
@@ A.1 Network Properties: Centrality and Clustering Metrics
We analyze the structural role of each agent in the directed follow-network $G=(V,E)$ with $N=|V|$ nodes by computing degree, betweenness, closeness, and eigenvector centralities, as well as the local clustering coefficient and global transitivity.
Degree centrality quantifies how many direct connections a node has relative to the maximum possible. For node $v$ with total degree $\deg(v)=k_{v}$ (in- plus out-degree), we define
@@
Closeness centrality reflects how near a node is to all others based on shortest-path distances $d(v,u)$. We set
@@
and the network-wide clustering is
@@
In these definitions, $d(v,u)$ denotes the shortest directed-path length from $v$ to $u$, $\sigma_{st}$ the total number of shortest paths from $s$ to $t$, and $\sigma_{st}(v)$ those passing through $v$. All sums run over $V$ unless noted otherwise.
In our simulated community of 161 agents engaging with 4,249 pieces of content, the resulting follow-network comprises 358 directed ties, indicating a modest level of connectivity driven by personas and memory-based decisions. An average degree centrality of 0.1105 shows that each agent follows roughly 11% of the population, while the low average betweenness centrality (0.0128) suggests that few agents act as indispensable bridges. With an average closeness centrality of 0.1367, most agents remain just a few steps apart, supporting rapid information flow, and the modest eigenvector centrality (0.0433) reveals that influence is fairly distributed rather than monopolized by a handful of hubs. Finally, a clustering coefficient of 0.205 and network transitivity of 0.189 point to moderate local grouping without excessive fragmentation—together painting a picture of a network that balances cohesion and reach in propagating popular content.
@@ A.2 Power-Law Distribution of User Popularity
First we define the popularity of users as a sum of the number of followers, number of likes, shares, and comments received. We collected the top 50 users and plotted their popularity (as measured by the sum of engagement received by them) from highest to lowest in Fig. 8. We observe a power-law distribution of user influence. We have $f(x)=120x^{-0.6}$ as the best-fitting power-law approximation of our sampled data, as shown in Fig. 5. With $\alpha=0.60$, our regression line has an $R^{2}=0.84$, suggesting a strong fit to our user engagement data and that our social system follows a typical power law distribution where a few users generate most of the engagement. Existing analysis on real-world social networks suggest that this power-law exponent usually ranges from 1.5-2.5 depending on the specific context (Muchnik et al., 2013; Bild et al., 2015). Our best-fit exponent is lower than these reported numbers, but it still illustrates a clear trend that a minority of users/content collect most of the engagement while the majority of them do not contribute nearly as much.
In the rest of this section, we explore potential reasons why this distribution emerges, and through a series of analyses leveraging our simulated environment, we reveal the unpredictability of influence or popularity in online social networks. More fundamentally, we argue that perhaps LLM-driven agents have a tendency to simply copy the decisions of agents who act before them. This results in the preferential attachment and as a natural consequence establishes the power-law distribution of engagement pattern. Such a pattern does not necessarily stem from anything else such as user’s profile details, or the content they post about. And even their own "inner reasoning" might not reveal their true decision-making, which invites further investigation on the authenticity of LLM agents’ self-expressed reasoning traces.

### v3

@@ Abstract
We present a novel, open-source social network simulation framework MOSAIC where generative language agents predict user behaviors such as liking, sharing, and flagging content. This simulation combines LLM agents with a directed social graph to analyze emergent deception behaviors and gain a better understanding of how users determine the veracity of online social content. By constructing user representations from diverse fine-grained actual user personas, our system enables multi-agent simulations that model content dissemination and engagement dynamics at scale. Within this framework, we evaluate three different content moderation strategies with simulated misinformation dissemination, and we find that they not only mitigate the spread of non-factual content but also increase user engagement. In addition, we analyze the trajectories of popular content in our simulations, and explore whether simulation agents’ articulated reasoning for their social interactions truly aligns with their collective engagement patterns. We open-source our simulation software to encourage further research within AI and social sciences: https://github.com/genglinliu/MOSAIC
@@ 1 Introduction
In 2024, OpenAI reported that its platform was already being misused by covert influence operations to generate synthetic content diffused over social media OpenAI (2024). These internet manipulators exploit the fact that social networks have become a fundamental part of modern life, shaping public discourse, influencing political opinions, and facilitating the rapid spread of unverified human- and AI-generated content Aichner et al. (2021); Orben et al. (2022); Cinelli et al. (2021). While traditional social science methods such as surveys and observational studies have provided insights into human behavior, they often struggle to capture large-scale, emergent online interactions Yu et al. (2021); Lorig et al. (2021). Agent-based modeling (ABM) provides distinct advantages over survey methods in social science research since it can simulate dynamic interactions over time, and support examination of hypothetical or counterfactual scenarios with repeatable and controllable conditions Bonabeau (2002); Epstein (1999).
@@
By bridging social science observations, game-theoretic modeling Acemoglu et al. (2023), and LLM-driven modeling, our work demonstrates the potential of generative agent simulations as a tool for studying large-scale online behaviors.
@@ 2 MOSAIC Social Network Simulation
We build a simulated social network environment inspired by platforms like X, [^1] https://x.com/, allowing AI-driven users to interact, post, and share content. The simulation includes a basic user class with attributes such as username, posts, followers, following, and reposts, mimicking the structure of real-world social media platforms. The network itself is defined by the follower-following relationships, creating a web of user interactions, represented by a directed graph $G=(N,E)$ where $N$ represents the set of user nodes, i.e., $N=\{n_{1},n_{2},\dots,n_{k}\}$, where $n_{i}$ is a user in the network. $E\subseteq N\times N$ represents the set of directed edges, i.e., $E=\{(n_{i},n_{j})\mid n_{i}\text{ follows }n_{j}\}$. Each edge $(n_{i},n_{j})$ signifies that user $n_{i}$ follows user $n_{j}$.
@@ 2.1 Simulation Flow
At each time step, news content is introduced based on predefined parameters, with agents dynamically responding to their feeds. Agents can optionally generate posts according to their own interests. However, during certain controlled experiments, we configure them to only engage through reactions such as liking, sharing, commenting, or reporting misinformation. We describe a more general action space and more details of their decision-making process in Appendix G. The visibility of posts evolves based on engagement metrics, simulating algorithmic amplification effects. If fact-checking is enabled, agents incorporate moderation signals, e.g., they are prompted to pay more attention to potentially falsified content or misinformation. We discuss the content moderation simulation in more depth in Section 3. Throughout a simulation run, the system tracks key statistics, including content reach, user influence, and misinformation spread. At the end of each simulation run, a post-hoc analysis is conducted to assess content diffusion dynamics, user engagement metrics, influence distribution, and the impact of fact-checking interventions. We also keep track of various network properties such as centrality [^2] The degree to which an user is central to the network, having outsized or undersized influence. and triadic closures, [^3] A common social phenomenon where users with a mutual connection are more likely to connect to each other. and perform homophily [^4] Another common social phenomenon where similar users are more likely to connect to each other over dissimilar ones. analysis to examine clustering patterns in user engagement (see Appendix A for an extended technical description of how we define and compute these network metrics). Our focus on these three properties is motivated by a large body of prior work in network theory and the social sciences Rapoport (1953); Granovetter (1973); McPherson et al. (2001); Abebe et al. (2022); Chang et al. (2025).
@@ 3.1 Setup
To collect factual news articles, we utilize a news aggregation API [^8] https://newsapi.org/ to retrieve articles published daily in legitimate sources from January 31 to February 28, 2025. The system queries the API for all available topics, prioritizing non-political popular articles in English. For each date in the range, we extract key information from the retrieved articles, including their title, description, and main content. Using the NewsAPI, we scraped a total of 2470 news articles from major media outlets.
@@
Our simulations involve agentic users interacting with news posts under four different fact-checking conditions: (1) No Fact-Checking, (2) Community-Based Fact-Checking, (3) Third-Party Fact-Checking with an independent LLM that uses its own parametric knowledge, and (4) Hybrid Fact-Checking, which integrates both community-based and third-party verification mechanisms. The simulations start with 50 agents and span over 40 time steps, with agents making interaction decisions based on the perceived veracity of posts and the presence of content moderation. At each new step, we randomly introduce up to 2 more agents into the environment to simulate the regular user growth of the social media platform. We analyze both the overall engagement with posts and the effectiveness of fact-checking strategies in suppressing misinformation.
@@
The action space of agents varies across different fact-checking conditions, reflecting different levels of scrutiny and intervention in their social media interactions. In the no fact-checking setting, agents interact freely with the feed, engaging with posts based solely on their interests and beliefs. They can like, share, comment, or ignore posts without any explicit instructions to assess the accuracy of the content. Under this setting, a post is considered “fact-checked" when it has at least one community note with a helpful rating. In the third-party fact-checking condition, the action space remains the same, but the environment implicitly assumes the presence of external fact-checkers who may influence the visibility or credibility of posts. However, the agents themselves do not perform any direct verification. In contrast, the community fact-checking setting expands the action space by allowing agents to add community notes to posts they deem misleading or in need of additional context, as well as rate existing community notes as either helpful or unhelpful. This introduces a participatory element, encouraging agents to contribute to a crowdsourced verification system. Finally, the hybrid fact-checking condition combines elements from both third-party and community-driven verification. Agents can engage with posts as in previous settings while also considering official fact-checks alongside community notes, contributing their own notes and rating those written by others. Across all conditions, agents must select from predefined valid actions, ensuring consistency in response formats. Additionally, when reasoning is enabled, agents are required to justify their interactions by providing a brief explanation for each chosen action, further enhancing the interpretability of their behavior.
@@ 3.2 False News Does Not Spread Faster than Real News with Simulation Agents
As shown in Fig. 4, factual news maintains consistently higher engagement levels than misinformation across all four settings. The gap is most pronounced under the Third-Party and Hybrid fact-checking conditions, where factual interactions climb steadily while misinfo remains low. Notably, in the No Fact-Checking scenario, false news still fails to gain a foothold, suggesting that these agents are inherently less likely to propagate unverified or misleading posts. We believe this behavior arises from two factors. First, the agents may rely on their internal confidence to guide sharing decisions, so low‐confidence (often false) content is passed over more often. And second, commercial LLMs have been post-trained with safety and helpfulness objectives. This training bias would make our agents less inclined to produce or circulate uncertain or harmful content.
@@ 3.3 Content Moderation Improves Both Fact-Checking and Engagement
The steadily widening gaps between the third-party/hybrid approaches indicate that certain content moderation generates compounding benefits for factual engagement over time. Among the simulated strategies, fact-checking by third parties proves most effective at building a healthy information ecosystem—an effect that aligns with human studies showing increased trust in unflagged content Pennycook et al. (2020). All curves in Fig. 4 represent metrics averaged across our three LLM agents (DeepSeek-V3, GPT-4o, and Claude-3.7-Sonnet). Interaction counts, including the number of shares, likes, comments, and notes, are aggregated at each time step to highlight overall platform dynamics under each moderation strategy. This high-level view illustrates consistent patterns in how content moderation shapes engagement. For a detailed, model-by-model breakdown and finer-grained metric analysis, please see Appendix J.
@@ 3.5 What Drives Users/Content Popularity?
Understanding what makes some content more popular is key to modeling engagement and intervention. In this standalone study, we simulate 161 agents over 4,249 posts of news and user-generated content, and let agents both react to and generate content based on personas, engagement cues, and memory-based decisions.
@@
Finally, we examine agents’ own reasoning traces to better understand engagement behaviors. Interestingly, agent reasoning traces show that sentiment and motivations do not always align with action types (e.g., positive sentiment in follows, likes, shares; quality or misinformation concerns in flags). As shown in Tab. 3, 13.5% of “flags” are coded as positive even though flag reasons are exclusively negative (information value, misinformation). The verbal reasoning for the “unfollow” actions is 20% positive, though it is intuitively a negative reaction towards other users. “ignore” omits any neutral reasoning and has a 77.8% positive rate in reasoning, which also makes their verbal explanations inconsistent with the actual action. Following these inconclusive analyses, we speculate that our recency-and-follow-based feed ranking creates preferential attachment: once followed, a user’s posts gain visibility, reinforcing their popularity. We provide further analyses in Appendix A.
@@ 5 Conclusion
Our findings are subject to several limitations, particularly in the scale of our experiments. First, the limited number of human participants, especially from minority demographic groups, restricts the statistical power of our conclusions. Expanding participant diversity would enable a more robust analysis of how alignment between real and simulated social interaction patterns varies across demographics.
Second, our content moderation experiments were conducted at a relatively small scale, which may have constrained the emergence of complex behaviors. Running these experiments at a larger scale could uncover additional dynamics not captured in the present study. Moreover, the simulation platform itself simplifies several aspects of real-world social media. For example, feed ranking is based primarily on recency and follow edges, rather than incorporating sophisticated recommender algorithms. This design choice was intentional to isolate the causal effects of content and moderation strategies, but it may limit the external validity of our findings. Future work could integrate recommender-driven dynamics while maintaining interpretability.
Finally, we observe a gap between agents’ explicit explanations for their actions and the collective reaction patterns that emerge in the system. The root causes of this misalignment remain unclear and warrant further investigation, potentially involving a deeper analysis of agent modeling assumptions or social influence mechanisms.
Ethical Considerations
This work raises several ethical considerations. First, our simulations incorporate misinformation narratives from the proprietary NewsGuard database. While these examples were necessary for experimental validity, they were used strictly for research purposes and are presented with appropriate warnings to prevent accidental amplification. We do not reproduce or promote harmful claims beyond what is needed for reproducibility.
Second, our human study with 204 participants was conducted under an approved protocol with informed consent. Participants were compensated fairly through Prolific. Sensitive demographic attributes were anonymized and used only in aggregate to generate agent personas. No personally identifying information is reported.
Third, our findings highlight that current LLMs exhibit safety alignment that reduces engagement with misinformation. While this property is beneficial in our study, it may not reflect real human susceptibility, and thus should not be interpreted as predictive of real-world behavior. We caution against overgeneralizing from simulation to deployment.
Finally, by releasing MOSAIC as an open-source framework, we aim to support responsible research in AI safety and computational social science. We encourage careful use of this tool, especially in studies that model or intervene in sensitive online behaviors, and emphasize that simulations should complement rather than replace empirical studies with human subjects.
@@ A.1 Network Properties: Centrality and Clustering Metrics
We analyze the structural role of each agent in the directed follow‐network $G=(V,E)$ with $N=|V|$ nodes by computing degree, betweenness, closeness, and eigenvector centralities, as well as the local clustering coefficient and global transitivity.
Degree centrality quantifies how many direct connections a node has relative to the maximum possible. For node $v$ with total degree $\deg(v)=k_{v}$ (in‐ plus out‐degree), we define
@@
Closeness centrality reflects how near a node is to all others based on shortest‐path distances $d(v,u)$. We set
@@
and the network‐wide clustering is
@@
In these definitions, $d(v,u)$ denotes the shortest directed‐path length from $v$ to $u$, $\sigma_{st}$ the total number of shortest paths from $s$ to $t$, and $\sigma_{st}(v)$ those passing through $v$. All sums run over $V$ unless noted otherwise.
In our simulated community of 161 agents engaging with 4,249 pieces of content, the resulting follow‐network comprises 358 directed ties, indicating a modest level of connectivity driven by personas and memory‐based decisions. An average degree centrality of 0.1105 shows that each agent follows roughly 11% of the population, while the low average betweenness centrality (0.0128) suggests that few agents act as indispensable bridges. With an average closeness centrality of 0.1367, most agents remain just a few steps apart, supporting rapid information flow, and the modest eigenvector centrality (0.0433) reveals that influence is fairly distributed rather than monopolized by a handful of hubs. Finally, a clustering coefficient of 0.205 and network transitivity of 0.189 point to moderate local grouping without excessive fragmentation—together painting a picture of a network that balances cohesion and reach in propagating popular content.
@@ A.2 Power-Law Distribution of User Popularity
First, we define the popularity of users as a sum of the number of followers, number of likes, shares, and comments received. We collected the top 50 users and plotted their popularity (as measured by the sum of engagement received by them) from highest to lowest in Fig. 8. We observe a power-law distribution of user influence. We have $f(x)=120x^{-0.6}$ as the best-fitting power-law approximation of our sampled data, as shown in Fig. 5. With $\alpha=0.60$, our regression line has an $R^{2}=0.84$, suggesting a strong fit to our user engagement data and that our social system follows a typical power law distribution where a few users generate most of the engagement. Existing analysis on real-world social networks suggests that this power-law exponent usually ranges from 1.5-2.5, depending on the specific context (Muchnik et al., 2013; Bild et al., 2015). Our best-fit exponent is lower than these reported numbers, but it still illustrates a clear trend that a minority of users/content collect most of the engagement, while the majority of them do not contribute nearly as much.
In the rest of this section, we explore potential reasons why this distribution emerges, and through a series of analyses leveraging our simulated environment, we reveal the unpredictability of influence or popularity in online social networks. More fundamentally, we argue that perhaps LLM-driven agents have a tendency to simply copy the decisions of agents who act before them. This results in the preferential attachment and, as a natural consequence, establishes the power-law distribution of engagement patterns. Such a pattern does not necessarily stem from anything else, such as the user’s profile details or the content they post about. And even their own "inner reasoning" might not reveal their true decision-making, which invites further investigation into the authenticity of LLM agents’ self-expressed reasoning traces.